A cylindrical rod has resistance R. If we triple its length and diameter, what is the resistance, in terms of R?


1
Expert's answer
2022-03-15T10:35:53-0400

Resistance of rod:

"R=\rho\cdot\frac{4l}{\pi\cdot d^2}"

Answer the resistance is R/3